# Dimensional crossover of class D real-space topological invariants

### v1

@@ I Introduction
The possible topological properties of non-interacting Hamiltonians are dictated by their spatial dimension, and discrete and crystal symmetries [1]. In particular, changing the dimension of a system while keeping its discrete symmetries unaltered generally yields a topological phase transition. Examples in condensed matter systems include transitions from a trivial insulator to a quantum spin Hall (QSH) phase in HgTe quantum wells [2, 3], and transitions from three-dimensional topological insulators with two-dimensional surface states to QSH states with one-dimensional helical states [4, 5, 6, 7].
Such topological transitions due to dimensional shifts have previously been characterized theoretically by analyzing edge states in finite-size systems, the topological invariants based on band representation [8, 9], and transport properties [10]. However, a real-space framework capable of illuminating how a system’s topology changes during a dimensional shift has traditionally been missing due to the dearth of local topological markers needed to probe the topology of systems lacking translational symmetry. Such a theory could be useful to guide the design of materials and next-generation devices, for example, to identify how thin a slab can be while maintaining its three-dimensional topological properties, or how thick a nanowire can be while maintaining its edge Majorana modes.
Figure 1: (a) Magnetic adatoms (filled circles) arranged in a circular island (a) and a chain (b) geometry, embedded in a two-dimensional superconductor whose sites are represented by empty circles. Panels (c) and (d) show the zero-energy local density of states (LDOS) for the circular Shiba island and chain, respectively, showing a Majorana mode at their boundary. Panels (e) and (f) show the LDOS as a function of energy at a bulk site (dashed black line) and at a site on the island’s and chain’s edges (solid orange line). The latter displays zero-energy peaks. The parameters used for the island and chain are $N=20$, $\mu=-4$, $\alpha=0.45$, $\Delta=2.0$, and $JS=4.8$
In this paper, we introduce a scheme to study topological phase transitions induced by the change in a system’s spatial dimensions via local topological markers. We will focus on the symmetry class D in the Atland-Zirnbauer classification [11]. Hamiltonians in this class have particle-hole symmetry, $\mathcal{C}$, and broken time-reversal and chiral symmetries. In one dimension (1D), the topological invariant corresponds to $\mathcal{Z}_{2}$, while in two dimensions (2D) to $\mathcal{Z}$ [12, 13]. Our aim is to study the evolution of the $\mathcal{Z}$ invariant into $\mathcal{Z}_{2}$ as one of the dimensions of the system is reduced. Some topological superconductors are examples of class D systems [14].
@@
The task of tracking the topological properties of the Shiba lattice across spatial dimensions is well suited to the spectral localizer [35, 36, 37], a real-space numeric K-theory approach that characterizes the local topological properties of finite-size systems, without requiring any notion of lattice periodicity. The spectral localizer approach has been employed to identify topological phenomena in gapless systems [38, 39, 40, 41], nonlinear materials [42], non-Hermitian systems [43, 44], photonic systems [45, 46, 47], quasicrystals [48, 49, 50], time-depedent systems [51, 52], disordered superconductors [53] and heterostructures [54].
We introduce two local topological markers derived from the spectral localizer, a Chern number, and a $\mathcal{Z}_{2}$ invariant that characterizes two and one-dimensional Shiba lattices embedded in two-dimensional superconductors. As these markers do not rely on projecting to an occupied subspace, they are endowed with a measure of the corresponding local topological protection related to the local spectral gap. The local nature of the markers permits us to track their changes and corresponding local protection as the dimension of the Shiba lattice changes from two to one dimension. The real-space topology in 2D Shiba lattices has been characterized using a global position-space invariant [55, 56], and $\mathcal{Z}_{2}$ local markers for nonchiral phases in odd dimensions have been recently introduced [57]. Still, so far, there is no local marker to diagnose the topology of 1D Shiba lattices embedded in a two-dimensional superconductor.
We will next introduce the specific model we consider, define the local topological markers, and present our results. We conclude this paper with a discussion of possible extensions of our work to other physical systems and other symmetry classes.
II Shiba lattice model Hamiltonian
A general mean-field tight-binding Hamiltonian for a Shiba lattice on a two-dimensional superconductor is given by [18, 56]
$\mathcal{H}$ $=-t\sum_{\mathbf{r},\mathbf{r}^{\prime},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}^{\prime}\sigma}-\mu\sum_{\mathbf{r},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma}+\Delta\sum_{\mathbf{r}}c^{\dagger}_{\mathbf{r}\uparrow}c^{\dagger}_{\mathbf{r}\downarrow}+h.c.$
@@ II Shiba lattice model Hamiltonian
where $c^{\dagger}_{\mathbf{r}\sigma}$ creates an electron at site $\mathbf{r}$ with spin $\sigma$, $\bm{\sigma}=\{\sigma_{x},\sigma_{y},\sigma_{z}\}$ are the Pauli matrices, $t$ is the hopping amplitude between nearest-neighbor lattice sites $\bm{\delta}$, $l=|\bm{\delta}|$, $\mu$ is the chemical potential, $\alpha$ is the spin-orbit coupling strength, and $\Delta$ is the superconducting s-wave pairing. The magnetic adatoms are located at positions $\mathbf{R}$ and are described as classical magnetic moments $\mathbf{S}$ aligned perpendicular to the superconductor lattice plane and coupled to the electron spin via the exchange coupling $J$.
In Fig. 1, we consider two magnetic adatom geometries: a circular Shiba island (a) and a chain (b), both embedded in a two-dimensional (square lattice) superconductor. We compute the local density of states (LDOS) at site $i$ and energy $E$,
@@
using the Chebyshev–Bogoliubov–de Gennes method [58], where $G^{\sigma\sigma}_{ii}(E)$ is the Green’s function. Panels (c) and (d) show the zero-energy LDOS, which shows Majorana modes localized at corresponding 1D and 0D edges for the circular island and chain, respectively. In panels (e) and (f), we show the LDOS as a function of energy at a bulk site and at an edge site of the magnetic adatom geometries. The former case shows the superconducting gap, while the latter shows the zero-energy peak. We considered here the representative parameters $N=20$, where $N$ is the size of the superconductor’s lattice, $\mu=-4$, $\alpha=0.45$, $\Delta=2$, and $JS=4.8$.
@@
Previously, the real-space topological properties of finite-size two-dimensional Shiba lattices have been characterized using real-space Chern numbers [55], including hybrid structures composed of networks of chains attached to islands [56]. Topological quantum numbers for one-dimensional chains based on the scattering matrix have been introduced [15]. A real-space topological invariant for one-dimensional time-quasiperiodic Majorana modes has been recently discussed in Ref. [52]. In contrast, the local topological marker we leverage here applies to one-dimensional Shiba chains embedded in a two-dimensional superconductor, such as the case in Fig. 1(d), relevant for experiments, and is endowed with a local measure of the topological protection.
In the next section, we will introduce the local topological marker that permits us to diagnose the topology of finite-size islands and chains of magnetic adatoms embedded in a two-dimensional superconductor and the crossover from 2D to 1D topology.
III Spectral localizer
@@
In this section, we introduce the local topological markers for 2D and 1D class D Hamiltonians. For a detailed tutorial on the spectral localizer and Clifford pseudospectrum (in the context of photonic systems), see Ref. [59]. Broadly, the spectral localizer framework seeks to classify material topology by first translating a system’s description into a single Hermitian matrix and then diagnose its topology via matrix homotopy, i.e., the study of which types of invertible matrices can be path-connected such that every element in the path shares the same type (e.g., Hermiticity) and remains invertible. As such, the spectral localizer framework does not rely on vector bundles, though it is provably equivalent in some cases [36].
@@
For a d-dimensional system with open boundaries described by the Hamiltonian $\mathcal{H}$ defined on a finite-dimensional Hilbert space, the spectral localizer is defined as [35]
$L_{\bm{\lambda}}$ $(\bm{X},\mathcal{H})=\sum_{j=1}^{d}\kappa(X_{j}-x_{j}\mathcal{I})\otimes\Gamma_{j}+\left(\mathcal{H}-E\mathcal{I}\right)\otimes\Gamma_{d+1},$ (3)
@@
where the matrices $\Gamma_{i}$ generate a Clifford representation, $\mathcal{I}$ is the identity matrix, $X_{j}$ are the position operators, and $\kappa>0$ is a scaling parameter with units of energy divided by length. Intuitively, $\kappa$ is chosen to balance the position and energy contributions to the spectrum of $L_{\bm{\lambda}}$. For bounded Hamiltonians, such as those describing tight-binding models, a proven range of validity exists [37]. Generally, for systems with gapped bulk spectra, this scaling parameter can be set by the bulk gap width $E_{\textrm{gap}}$ and the smallest length of the finite system $L$ as $\kappa\sim E_{\textrm{gap}}/L$. In comparison with other theories of local topological markers [60, 61], setting $\kappa$ is similar to the choice of integration radius used by those markers. Finally, $\bm{\lambda}=(\bm{x},E)$ is the point in position-energy space at which we are probing the local topological properties.
@@
The spectral localizer $L_{\bm{\lambda}}(\bm{X},\mathcal{H})$ encodes both the local topological invariant and a measure of its local protection. In particular, as this framework’s local markers are all rooted in matrix homotopy arguments, the spectral localizer locally quantifies topological protection by how close a system’s $L_{\bm{\lambda}}(\bm{X},\mathcal{H})$ is to being non-invertible at a chosen $\bm{\lambda}$. In other words, the local gap
$\mu_{\bm{\lambda}}(\bm{X},\mathcal{H})=\text{min}\left\{\left|\Sigma\left(L_{\bm{\lambda}}(\bm{X},\mathcal{H})\right)\right|\right\},$ (4)
@@
In the following sections, we specialize the spectral localizer for the 2D and 1D cases, and define in each case the property that characterizes whether the Hamiltonian and position operators of a class D system can be continued to an atomic limit, i.e., a set of commuting matrices, or not, without breaking parity and while maintaining a non-zero local gap.
Figure 2: (a) Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|$ and (b) 2D localizer gap as a function of position for the 2D island geometry shown in Fig. 1(a), for the same parameters. The spectral localizer calculations use $\kappa=0.05(t/l)$.
III.1 2D invariant
$L^{(2D)}_{\bm{\lambda}=(x,y,E)}$ $(X,Y,\mathcal{H})=\left(\mathcal{H}-E\mathcal{I}\right)\otimes\sigma_{z}$
$+\kappa\left[(X-x\mathcal{I})\otimes\sigma_{x}+(Y-y\mathcal{I})\otimes\sigma_{y}\right].$ (5)
For a Hamiltonian in class D, the local topological marker at the energy-position point $\bm{\lambda}=(x,y,E)$ corresponds to
$C_{\bm{\lambda}}^{\textrm{L}}=\frac{1}{2}\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,\mathcal{H}\right)\},$ (6)
where $\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,\mathcal{H}\right)\}$, corresponds to the signature of the spectral localizer, the difference between the number of positive and negative eigenvalues. When $\mu_{\bm{\lambda}}$ is nonzero, $C_{\bm{\lambda}}^{\textrm{L}}$ takes a well-defined integer value. In Fig. 2, panel (a), we plot $|C_{\bm{\lambda}}^{\textrm{L}}|$ as a function of position for the Shiba lattice in Fig. 1 (a). In panel (b), we show the corresponding local gap, $\mu_{\bm{\lambda}}$.
Figure 3: Topological phase diagram as a function of the chemical potential $\mu$ and $J$, for $N=20$, and $\alpha=0.45$. Panel (a) shows the case $\Delta=1.2$, while panel (b) $\Delta=0.4$. The color code indicates the local topological marker at the center of a circular island, $|C_{\bm{\lambda}}^{\textrm{L}}|$, and at the center of a chain, $z_{2,\bm{\lambda}}$, using the same parameters for both geometries. Topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. The spectral localizer calculations use $\kappa=0.05(t/l)$.
Figure 4: (a)-(e): Arrangements of magnetic adatoms interpolating from the two-dimensional circular island to the one-dimensional chain geometries. (f)-(j): 1D (black solid line) and 2D (dashed orange lines) localizer spectrum $\Sigma(L_{\bm{\lambda}})$ as a function of position $x$ for $y=N/2$. (k)-(o): Local Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|$ (orange squares) and 2D localizer gap $\text{min}\{|\Sigma(L^{(2D)}_{\bm{\lambda}})|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. (p)-(t): Local $z_{2,\bm{\lambda}}$ invariant (black circles) and 1D localizer gap $\text{min}\{\left|\Sigma(L^{(1D)}_{\bm{\lambda}})\right|\}$ (solid gray line) as a function of position $x$ for $y=N/2$. The parameters used are the same as in Fig. (1). The spectral localizer calculations use $\kappa=0.05(t/l)$.
III.2 1D invariant
@@
$L^{(1D)}_{\bm{\lambda}=(x,E)}(X,\mathcal{H})=\left(\mathcal{H}-E\mathcal{I}\right)\otimes\sigma_{y}+\kappa(X-x\mathcal{I})\otimes\sigma_{x},$ (7)
@@
where $\mathcal{H}$ is also given by Eq. (1). Note that the choice of $\mathcal{H}$ need not change bewteen using $L^{(2D)}_{\bm{\lambda}}$ and $L^{(1D)}_{\bm{\lambda}}$. Instead, the dimensional reduction is performed through the omission of the second position operator $Y$ in the definition of $L^{(1D)}_{\bm{\lambda}}$. As such, $L^{(1D)}_{\bm{\lambda}}$ can be equally applied to 1D chains as well as effective 1D chains embedded in 2D materials (or any material of arbitrarily higher dimension). Similarly, one can also define $L^{(1D)}_{\bm{\lambda}}$ in terms of other position operators such as $Y$ or $X+Y$. However, for concreteness, we arrange our Shiba chains along the x-direction in the 2D superconductor.

### v2

@@ I Introduction
The possible topological properties of non-interacting Hamiltonians are dictated by their spatial dimension, and discrete and crystal symmetries [1]. In particular, changing the dimension of a system while keeping its discrete symmetries unaltered generally yields a topological phase transition, following the Altland-Zirnbauer classes [2, 3, 4, 5]. Examples in condensed matter systems include transitions from a trivial insulator to a quantum spin Hall (QSH) phase in HgTe quantum wells [6, 7], and transitions from three-dimensional topological insulators with two-dimensional surface states to QSH states with one-dimensional helical states [8, 9, 10, 11].
Such topological transitions due to dimensional shifts have been previously theoretically characterized by analyzing edge states in finite-size systems, the topological invariants based on band representations [12, 13], and the transport properties of a material [14]. However, a real-space framework capable of illuminating how a system’s topology changes during a dimensional shift has traditionally been missing due to the dearth of local topological markers needed to probe the topology of systems lacking translational symmetry across all of the possible symmetry classes. Nevertheless, such a theory could be useful to guide the design of materials and next-generation devices; for example, to identify how thin a slab can be while maintaining its three-dimensional topological properties, or how thick a nanowire can be while maintaining its end Majorana modes.
Figure 1: Topological phenomena in 1D and 2D Shiba lattices. a,b Magnetic adatoms (filled circles) arranged in a circular island (a) and a chain (b) geometry, embedded in a two-dimensional superconductor whose sites are represented by empty circles. c,d Zero-energy local density of states (LDOS) for the circular Shiba island and chain, respectively, showing a Majorana mode at their boundary. e,f LDOS as a function of energy at a bulk site (dashed black line) and at a site on the island’s and chain’s edges (solid orange line). The latter displays zero-energy peaks. The parameters used for the island and chain are $N=20$, where $N$ is the size of the superconductor’s lattice, chemical potential $\mu=-4t$, spin-orbit coupling strength $\alpha=0.45t/l$, superconducting s-wave pairing $\Delta=2.0t$, and magnetic moment times exchange coupling $JS=4.8t$.
In this paper, we introduce a scheme to study topological phase transitions induced by the change in a system’s spatial dimensions via local topological markers. We will focus on the symmetry class D in the Altland-Zirnbauer classification [2, 3, 4, 5], a class that contains some topological superconductors [15], and which exhibits non-trivial invariants in both 1D and 2D. Specifically, Hamiltonians in this class have particle-hole symmetry $\mathcal{C}$, and broken time-reversal and chiral symmetries. In one dimension (1D), the topological invariant corresponds to $\mathbb{Z}_{2}$, while in two dimensions (2D) to $\mathbb{Z}$. Our aim is to study the evolution of the $\mathbb{Z}$ invariant into $\mathbb{Z}_{2}$ as one of the dimensions of the target material embedded in an insulating environment is reduced.
@@
The task of tracking the topological properties of the Shiba lattice across spatial dimensions is well suited to the spectral localizer framework [36, 37, 38], a real-space numeric $K$-theory approach that characterizes the local topological properties of finite-size systems, without requiring any notion of lattice periodicity. The spectral localizer approach has been employed to identify topological phenomena in gapless systems [39, 40, 41, 42], nonlinear materials [43], non-Hermitian systems [44, 45, 46, 47, 48], photonic systems [49, 50, 51], quasicrystals [52, 53, 54], time-dependent systems [55, 56], disordered superconductors [57] and heterostructures [58]. While most of these studies have looked at 1D and 2D systems, the framework has also been applied to a 3D system [36] and a 4D system [59].
For the present study, we utilize two local topological markers derived from the spectral localizer, a Chern number, and a $\mathbb{Z}_{2}$ invariant that characterizes two- and one-dimensional Shiba lattices embedded in two-dimensional superconductors. Additionally, we also considered the one-to-zero-dimensional crossover in a BDI class model, showing the generality of our approach (See Supplementary Note 1). As this framework does not rely on projecting to an occupied subspace, it is endowed with a measure of the corresponding local topological protection related to the local spectral gap [60]. The local nature of the markers permits us to track their changes and corresponding local protection as the dimension of the Shiba lattice changes from two to one dimension. The real-space topology in 2D Shiba lattices has been characterized using a global position-space invariant [61, 62], and $\mathbb{Z}_{2}$ local markers for nonchiral phases in odd dimensions have been recently introduced [63]. However, these approaches cannot diagnose the topology of 1D Shiba lattices embedded in a two-dimensional superconductor.
II Results
II.1 Shiba lattice model Hamiltonian
A mean-field tight-binding Hamiltonian describing Shiba lattices, systems consisting of a set of magnetic atoms placed on the surface of a superconductor, is given by [19, 62]
$H$ $=-t\sum_{\mathbf{r},\bm{\delta},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}+\bm{\delta}\sigma}-\mu\sum_{\mathbf{r},\sigma}c^{\dagger}_{\mathbf{r}\sigma}c_{\mathbf{r}\sigma}+\Delta\sum_{\mathbf{r}}c^{\dagger}_{\mathbf{r}\uparrow}c^{\dagger}_{\mathbf{r}\downarrow}+h.c.$
@@ II.1 Shiba lattice model Hamiltonian
where $c^{\dagger}_{\mathbf{r}\sigma}$ ($c_{\mathbf{r}\sigma}$) creates (annihilates) an electron at site $\mathbf{r}$ with spin $\sigma$, $\bm{\sigma}=\{\sigma_{x},\sigma_{y},\sigma_{z}\}$ are the Pauli matrices, $t$ is the hopping amplitude between nearest-neighbor lattice sites separated by the displacement vector $\bm{\delta}$ such that the distance between nearest-neighbor sites is $l=|\bm{\delta}|$, $\mu$ is the chemical potential, $\alpha$ is the spin-orbit coupling strength, and $\Delta$ is the superconducting s-wave pairing. Specifically, we consider a square lattice superconductor. The magnetic adatoms located at positions $\mathbf{R}$ are assumed single-orbital, and are described as classical magnetic moments $\mathbf{S}$ aligned perpendicular to the superconductor lattice plane and coupled to the electron spin via the exchange coupling $J$.
Shiba lattices possess particle-hole symmetry $\mathcal{C}^{2}=I$, where $I$ is the identity, such that they are in class D of the Altland-Zirnbauer classification and can thus exhibit non-trivial topology in physically realizable dimensions. To illustrate the topological behavior of Shiba lattices, we consider two magnetic adatom geometries in Fig. 1: a circular Shiba island (a) and a chain (b), both embedded in a two-dimensional square-lattice superconductor. We compute the local density of states (LDOS) at site $i$ and energy $E$,
@@
using the Chebyshev–Bogoliubov–de Gennes method [64], where $G^{\sigma\sigma}_{ii}(E)$ is the Green’s function, and $\sigma$ corresponds to the spin. We include an artificial broadening of the Green’s function’s features at zero temperature [see Methods]. Majorana modes localized at the corresponding 1D and 0D edges of the circular island and chain are numerically observed in their respective zero-energy LDOS, see Fig. 1(c) and (d). Moreover, we compare the LDOS as a function of energy at bulk and edge sites of the magnetic adatom geometries in Fig. 1(e) and (f); the former case shows the superconducting gap, while the latter exhibits a zero-energy peak corresponding to the interface-localized states.
@@
Previously, topological properties of finite-sized two-dimensional Shiba lattices have been classified in real-space using Chern markers [61], including hybrid structures composed of networks of chains attached to islands [62]. Topological quantum numbers for one-dimensional chains based on the scattering matrix have been introduced [16]. Similarly, scattering theory has been recently applied to diagnose intrinsic higher order topology protected by spatial symmetries in finite samples [65]. A real-space topological invariant for one-dimensional time-quasiperiodic Majorana modes has been discussed in Ref. [56].
II.2 Spectral localizer
@@
Here, we use the spectral localizer framework to study the experimentally relevant geometry of a one-dimensional Shiba chains embedded in a two-dimensional superconductor, such as the case in Fig. 1(d). In general, this framework has two key features that make it amenable to this task: first, it possesses local topological markers for both 1D and 2D class D systems, and second, it is endowed with a local measure of the associated topological protection. Physically, the spectral localizer framework assesses whether a system can be locally continued to an atomic limit while preserving any local spectral gap and (relevant) symmetries, or if there is an obstruction to doing so. Mathematically, the spectral localizer framework classifies material topology by first translating a system’s description into a single Hermitian matrix and then diagnoses its topology via matrix homotopy, i.e., the study of which types of invertible matrices can be path-connected such that every matrix element in the path shares the same properties (e.g., Hermiticity) and remains invertible. As such, the spectral localizer framework does not rely on vector bundles, though it is provably equivalent in some cases [37, 66]. For a detailed tutorial on the spectral localizer and Clifford pseudospectrum, see Ref. [67].
@@
For a $d$-dimensional system with open boundaries described by the Hamiltonian $H$ defined on a finite-dimensional Hilbert space, the spectral localizer is defined as [36]
$L_{\bm{\lambda}}$ $(\bm{X},H)=\sum_{j=1}^{d}\kappa(X_{j}-x_{j}I)\otimes\Gamma_{j}+\left(H-EI\right)\otimes\Gamma_{d+1},$ (3)
@@
where the matrices $\Gamma_{i}$ generate an irreducible Clifford representation, $X_{j}$ are the position operators, $\bm{\lambda}=(\mathbf{x},E)$ is the point in position-energy space at which we are probing the local topological properties, and $\kappa>0$ is a scaling parameter with units of energy divided by length. Intuitively, the spectral localizer can be understood as combining the eigenvalue equations of the position operators with the Hamiltonian into a single matrix using a Clifford representation to preserve the independent information contained in these different operators. Thus, $\kappa$ must be chosen to balance the relative contributions of $X_{j}$ and $H$ to the spectrum of $L_{\bm{\lambda}}$; choosing $\kappa$ to be too large or too small results in the information carried by $L_{\bm{\lambda}}$ to effectively duplicate the information already present in $X_{j}$ or $H$, respectively. For bounded Hamiltonians, such as those describing tight-binding models, a proven range of validity for $\kappa$ exists where the spectral localizer provides information beyond what is immediately available in $H$ and $\bm{X}$ [38, 60, 66]. Generally, for systems with gapped bulk spectra, this scaling parameter can be set by the bulk gap width $E_{\textrm{gap}}$ and the smallest length of the finite system $L$ as $\kappa\sim E_{\textrm{gap}}/L$. In comparison with other theories of local topological markers [68, 69], setting $\kappa$ is similar to the choice of integration radius used by those markers.
@@
The spectral localizer $L_{\bm{\lambda}}(\bm{X},H)$ encodes both the local topological invariant and a measure of its local protection. In particular, as this framework’s local markers are all rooted in matrix homotopy arguments, the spectral localizer quantifies the strength of the local topological protection by how close a system’s $L_{\bm{\lambda}}(\bm{X},H)$ is to being non-invertible at a chosen $\bm{\lambda}$. More precisely, the localizer gap is defined as
$\mu_{\bm{\lambda}}(\bm{X},H)=\text{min}\left\{\left|\Sigma\left(L_{\bm{\lambda}}(\bm{X},H)\right)\right|\right\},$ (4)
where $\Sigma\left(L_{\bm{\lambda}}(\bm{X},H)\right)$ corresponds to the spectrum of the spectral localizer. In other words, this localizer gap is the minimum distance an eigenvalue must move for $L_{\bm{\lambda}}$ to become non-invertible. As the spectral localizer is Hermitian (for Hermitian $H$), any perturbation $\delta L$ to $L_{\bm{\lambda}}$, either through changing the choice of $\bm{\lambda}$ or perturbing the underlying system $H\rightarrow H+\delta H$, cannot yield a non-invertible $L_{\bm{\lambda}}+\delta L$ so long as $\|\delta L\|<\mu_{\bm{\lambda}}$. Here, $\|\cdot\|$ is the $L_{2}$ matrix norm, i.e., its largest singular value.
The localizer gap can also be viewed as a quantity similar to the local density of states. As the spectral localizer is a combination of generally non-commuting operators, $\bm{\lambda}=(\mathbf{x},E)$ can be understood as a guess at a location in position and energy space where the operators might exhibit an approximate joint eigenvector $|\phi\rangle$. If $\mu_{\bm{\lambda}}$ is small relative to quantities like $\sqrt{\|[H,\kappa X_{j}]\|}$, the system is guaranteed to exhibit a state that satisfies $X_{j}|\phi\rangle\approx x_{j}|\phi\rangle$ and $H|\phi\rangle\approx E|\phi\rangle$ [70]. Thus, altogether, large values of $\mu_{\bm{\lambda}}$ indicate locations where the local topology is robust, while small values of $\mu_{\bm{\lambda}}$ indicate the presence of approximate states in the underlying system.
Figure 2: Application of the spectral localizer framework to the Shiba island. a,b Chern number $|C_{\bm{\lambda}}^{\textrm{L}}|$ (a) and 2D localizer gap (b) as a function of position for the 2D island geometry shown in Fig. 1a, for the same parameters and at $E/t=0$. The spectral localizer calculations use $\kappa=0.05(t/l)$.
II.3 2D class D local invariant
While the definitions of the spectral localizer and the associated localizer gap can be defined for arbitrary dimensions, the associated local markers for classifying material topology require specializing to both the specific system dimensionality and symmetry class being considered. This demand for specificity can both be viewed as a physical consequence of the Altland-Zirnbauer classification [3, 4, 5], as well as a mathematical consequence of changing dimensions and symmetries altering the structure of $L_{\bm{\lambda}}(\bm{X},H)$. For example, such alterations may transform a spectral localizer that is symmetric $L_{\bm{\lambda}}^{\top}=L_{\bm{\lambda}}$ into one that no longer possesses this property, fundamentally changing what matrix homotopy arguments can be applied and thus changing the structure, and possible existence, of the local markers.
$L^{(2D)}_{\bm{\lambda}=(x,y,E)}$ $(X,Y,H)=\left(H-EI\right)\otimes\sigma_{z}$
$+\kappa\left[(X-xI)\otimes\sigma_{x}+(Y-yI)\otimes\sigma_{y}\right].$ (5)
For a Hamiltonian in class D, the local topological marker at the energy-position point $\bm{\lambda}=(x,y,E)$ is [36]
$C_{\bm{\lambda}}^{\textrm{L}}=\frac{1}{2}\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,H\right)\},$ (6)
where $\text{sig}\{L^{(2D)}_{\bm{\lambda}}\left(X,Y,H\right)\}$ corresponds to the signature of the spectral localizer, the difference between the number of positive and negative eigenvalues. Thus, when $\mu_{\bm{\lambda}}$ is nonzero, $C_{\bm{\lambda}}^{\textrm{L}}$ takes a well-defined integer value. Note that $C_{\bm{\lambda}}^{\textrm{L}}$ can only change its value across $\bm{\lambda}$ where $\mu_{\bm{\lambda}}=0$, as the eigenvalues of Hermitian operators must move continuously.
The overall picture of the spectral localizer framework is shown in Fig. 2, where the local marker $|C_{\bm{\lambda}}^{\textrm{L}}|$ and localizer gap $\mu_{\bm{\lambda}}$ are both shown over the spatial extent of the full 2D system. In the lattice’s center the local marker reveals non-trivial topology and the localizer gap is large, indicating the robustness of the topological phase. As the choice of position is varied towards the system’s edge, the localizer gap closes around the boundary of the distribution of magnetic adatoms, beyond which the local index changes its value to reveal trivial topology. Finally, at the heterostructure’s boundary, the closing of the localizer gap coincides with the chiral interface-localized states, see Fig. 1c. Thus, bulk-boundary correspondence manifests in the spectral localizer framework as the local index cannot change without crossing through a location in position-energy space where $\mu_{\bm{\lambda}}=0$, guaranteeing the appearance of a system state at that approximate energy and position.
Figure 3: Dimensional crossover topological phase diagram in the Shiba lattice. The phase diagram is shown as a function of the chemical potential $\mu/t$ and the product of the magnetic moment and exchange coupling $JS/t$, for $N=20$, and spin-orbit coupling strength $\alpha=0.45t/l$. The phase diagrams contain $31\times 31$ points. a-c Phase diagram (a), 1D spectral localizer gap for the chain (b), and 2D spectral localizer gap for the chain (c) for $\Delta=1.2t$ at $E=0$. The color code indicates the 2D local marker $|C_{\bm{\lambda}}^{\textrm{L}}|$ computed at the center of a circular island, and the 1D local marker $z_{2,x}$ computed at the center of a chain using the same parameters and only changing the geometry. The localizer gaps are similarly computed at the center of the 2D island and 1D chain. d-f Similar to a-c, except for $\Delta=0.4t$. Topological transitions not captured in the diagram can occur as the geometry changes from 2D to 1D. The spectral localizer calculations use $\kappa=0.05(t/l)$.
@@
[…]
II.4 1D class D local invariant
@@
$L^{(1D)}_{\bm{\lambda}=(x,E)}(X,H)=\left(H-EI\right)\otimes\sigma_{y}+\kappa(X-xI)\otimes\sigma_{x},$ (7)
@@
where $H$ is also given by Eq. (1). Here, the choice of $H$ need not change between using $L^{(2D)}_{\bm{\lambda}}$ and $L^{(1D)}_{\bm{\lambda}}$, that is, the same $H$ could be used in both Eqs. (5) and (7) to determine different physical properties. Instead, to classify topological phenomena associated with a lower physical dimension we use dimensional reduction, which in the spectral localizer framework is performed through the omission of (in this case) the second position operator $Y$ in the definition of $L^{(1D)}_{\bm{\lambda}}$. Intuitively, the absence of a position operator means that sites with different $y$ coordinates but the same $x$ coordinate are being treated mathematically as analogous to different orbitals of a single site with the same $x$ coordinate, a process that could be considered as similar to the reverse of using a synthetic dimension [71] where different degrees of freedom on a single spatial site are considered to constitute an additional fictitious spatial dimension. As such, $L^{(1D)}_{\bm{\lambda}}$ can be equally applied to 1D chains as well as effective 1D chains embedded in 2D materials (or any material of arbitrarily higher dimension). Similarly, one can also define $L^{(1D)}_{\bm{\lambda}}$ in terms of other position operators such as $Y$ or $X+Y$. However, for concreteness, we arrange our Shiba chains along the $x$-direction in the 2D superconductor as shown in Fig. 1b.